OpenStudy (anonymous):

Evaluate the following integral: integral (x^3)/square root (x^2 + 16) from 0 to 3

hartnn (hartnn):

did you try u substitution ? u = x^2+16 [so, x^2 = u-16] du =... ?

hartnn (hartnn):

yes/no/trying ?

OpenStudy (anonymous):

trying for sure! :) Mmm... not quite sure about the du part.. can u explain that to me please

hartnn (hartnn):

u= x^2+16 differentiate this w.r.t x du/ dx =...?

hartnn (hartnn):

the catch here is to split the numerator's x^3 dx \(x^3dx = x^2 (x dx)\)

OpenStudy (anonymous):

hmm.. im still a bit confused... so when i find du i get 2x dx then im stuck idk where to go from there :(

hartnn (hartnn):

yes, so du = 2x dx, right ? so, x dx = du/2 got this ?

OpenStudy (anonymous):

yup

hartnn (hartnn):

Numerator = x^3 dx = x^2 (x dx) = (u-16) (du/2) and this ?

hartnn (hartnn):

any doubts in numerator ? denominator is just sqrt u, isn't it

OpenStudy (anonymous):

ahhh okie dokie got it

hartnn (hartnn):

good! go you can solve further ?

hartnn (hartnn):

*so

OpenStudy (anonymous):

yup i think i'll be ok :) thank you! If I still get stuck on smthg along the way can i still bug you for help? :P sorry kinda new to this..

hartnn (hartnn):

ofcourse! no problem at all :) since you're new here, WELCOME to OpenStudy!! :)